OpenStudy (anonymous):

a triangle has an area of 24 sq.m. the base 2 m longer than the height. what is the length of the base of the triangle?

OpenStudy (anonymous):

Base= 8 m

OpenStudy (anonymous):

\[A=24\] \[A=\frac{ 1 }{ 2 }(B)(H)\] ( where B= H+2) \[24=\frac{ 1 }{ 2}( h+2)(h)\] \[48 = h^2+2h\] \[h^2+2h-49=0\] \[(h+8)(h-6)=0\] H= -8 , 6 ( -8 is not possible) Solving for B B= h+2 = 6+2 =8

OpenStudy (anonymous):

thnx